OpenStudy (anonymous):

Triangle ABC has side a = 12, b = 16, and c = 19. Find the measure of angle C to the nearest whole number.

OpenStudy (asnaseer):

Use the cosine rule here. If you are not familiar with this rule then this site might help explain it: http://www.y-maths.co.uk/trisincosrl.htm

OpenStudy (anonymous):

Show Steps! Plz

OpenStudy (asnaseer):

have you read the details on the link I posted?

OpenStudy (anonymous):

Yess

OpenStudy (asnaseer):

so how far did you get before you got stuck?

OpenStudy (anonymous):

Im not sure how to get angel A. I got 39 but it saids is not the answer i dont know what i did wrong

OpenStudy (asnaseer):

please show your steps so that I can spot where you went wrong

OpenStudy (anonymous):

\[Cos A= 12^{2}-16^{2}-19^{2} \div(-2(16)(19))\] \[Cos A=(14-256-361)/-608\] \[Cos A= 0.7780 =39\]

OpenStudy (asnaseer):

the cosine rule states, for angle A:\[a^2=b^2+c^2-2bc\cos(A)\]therefore:\[\cos(A)=\frac{b^2+c^2-a^2}{2bc}\]

OpenStudy (asnaseer):

however, you state you want angle C in your question, in which case it should be:\[\cos(C)=\frac{a^2+b^2-c^2}{2ab}\]

OpenStudy (anonymous):

So I dont need to find angle A to get to C?

OpenStudy (asnaseer):

why find angle A when you can find angle C directly using the cosine rule?

OpenStudy (anonymous):

I Though I Had To . Thank You I Got The Answer Now (:

OpenStudy (asnaseer):

ok - glad you were able to work it out in the end :)